Solve the given integral equation or integro-differential equation for yt). t yt) + (t-vy(v) dv=51, y(0)=0 0 y(t) =
Answer #1

Solution → we have integral equation, y la +1 (t- you dve eta y lo) o = 0 Diffesentiateig co wirot towe have we used Rule. y=0 Now, yo) From a From (2) y20)=0 4110) =5 llut t=0 inw) (put t=0 is 2) we haue y ty (*7 - EO yio =0 yeo) =0 -5 ylo) Aunil:D Now, (orih = 0 .C +(2 2 have -(6) +62 + 13 C3 Differentiate yºlt zwe g (t) cient -12 et/2 Los( Bt) sur ( t) + { Czetla co:. Cias 3 and C3 s 3 t/2 te yet) = Set cos 13 It + -5 3 3 2 t 품 9 sein / 3 (G 3 of gwen integral is thre Solutwin equations

PLEASE GIVE THIS A RATING IF THIS WAS HELPFUL FOR YOU